Considering the volume of a spherical shell as an increment of volume of sphere, find approximately the volume of a spherical shell whose outer diameter is 8 inches and whose thickness is 1/16 inch.


1
Expert's answer
2021-01-20T02:53:43-0500

Outer sphere volume: (4/3)*"\pi" *R3, R - outer diameter.

Inner sphere volume: (4/3)*"\pi" *r3, r - inner diameter.

The difference between the volumes will give the shell volume:

4/3*"\pi" *(R3-r3) = 4/3*"\pi"*(R-r)*(R2+R*r + r2) -> (R-r) = 1/16 inch -> 4/3*"\pi"*(1/16)*(R2+R*r+r2) =

= (1/3*1/4)*"\pi" *(82 + 8*"(8-\frac{1}{16})" + (8-"\\frac{1}{16}\n\n\u200b" )2 ) = 1/12 * (82 + 82 - "\\frac{8}{16}\n\u200b" + (82 - 2*8*"\\frac{1}{16}\n\n\u200b" + "\\frac{1}{256}\n\n\u200b"))*"\pi" =

= "\\frac{1}{12}\n\n\u200b" * (3*82 - 1.5 + "\\frac{1}{256}\n\n\u200b")*"\pi" = ("\\frac{1}{4}\n\n\u200b"*64 - 1.5 +"\\frac{1}{3072}\n\n\u200b") * "\pi" = (14.5 + 0.0003) * "\pi" "\approx" 14.5003 * 3.14159 "\approx" 45.554 cubic inches (inch3).

Answer: 45.554 inch3.